1/1

Mèng Hàorán,  Mèng Hàorán,
Sommerabend  The glow on western mountains
tradukita de Manfred Hausmann  tradukita de N. N. 81

Der Berge goldnes Licht vergeht im abendlichen Schattenreich. Mit zitterndem Geblink entsteht das Bild des Mondes auf dem Teich. Gelösten Haares liege ich auf der Terrasse, und mein Sinn, mein Leben gibt der Stille sich, gibt sich dem kühlen Mondlicht hin. Ein unbestimmtes Wehen streicht, nach Steinklee duftend, um das Haus. Der Bambusschatten regt sich leicht und rauscht vom Tropfenfall des Taus.

The glow on western mountains quickly sets, The moon is climbing over the eastern lake. My hair loose, I enjoy the evening cool, I lie in peace before the open window. The wind spreads lotus scent all through the air, The sound of dripping bamboo dew is clear. Although I’d like to fetch my qin and play, To my regret, there is no-one to hear. So touched by this, I think of my old friend, Throughout the night, I’m troubled by my dreams.

Traduko de la ĉina poemo “Shān guāng” de Mèng Hàorán (∗689 – †740) en la Germanan de Manfred Hausmann (∗1898-09-10 – †1986-08-06).

Arg-995-2007 (2014-04-28 16:51:56)

La germana poeto Manfred Hausmann verkis tiun ĉi germanigon de la poemo laŭ laŭvorta germanigo de la origina ĉina poemo.

Traduko de la ĉina poemo “Shān guāng” de Mèng Hàorán (∗689 – †740) en la Anglan de N. N. 81.

Arg-995-2010 (2014-04-20 00:35:06)

Tiu ĉi laŭsenca angligo troviĝas en http: // www. chinese-poems. com/ m6. html .

http://www.poezio.net/version?poem-id=995&version-id=,2007,2010